Operations Report – Mar 31st

Crew 208 Operations Report 31-MAR-2019

SOL: 3

Name of person filing report: Amanda Manget

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and alternator need replacement on the generator. A contractor has been arranged to make these repairs within a few days.

Generator (hours run): 12.5

Solar – SOC 66% @ 1900 (before David turned on generator for the night)

Diesel Reading – 47%

Propane Reading – 54%

Ethanol Free Gasoline – 2 gallons

Water (auxiliary tank) – 0 gallons (Note: Not in Use)

Water (static tank) – 194 gallons

Auxiliary to Static tank transfer – No

Gallons transferred: 0

Water in GreenHab – 217 gallons

Static to Loft Pump used – Yes

Water Meter: 014225.5 gallons @ 1822

Toilet tank emptied: No

Deimos rover used: No, not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Spirit rover used: Yes

Hours: 85.3

Beginning charge: 100%

Ending charge: 68%

Currently charging: Yes

Opportunity rover used: Yes

Hours: 55.7

Beginning charge: 100%

Ending charge: 78%

Currently charging: No

Curiosity rover used: Yes

Hours: 85.9

Beginning charge: 100%

Ending charge: 69%

Currently charging: Yes

Notes on rovers:

Dean, David, Erika and I took the Opportunity and Curiosity rovers out to the Candor Chasma (out of sim). On Cactus Road 1104 (at 4252000 and 521000), a high temperature battery warning came on the Opportunity rover while driving straight, and it stopped suddenly. We tried to communicate with David but there was no signal. We waited approximately 20 minutes to wait for it to cool down, and we took Curiosity to a higher point to try to establish contact again, which we could not. Then we checked Opportunity and the warning was no longer there so we started on the path again. About 5 minutes later it stopped again. Then we took Curiosity to a high point again and were able to establish contact with Julielynn who was able to reach David. David then came on an ATV to tow Opportunity back. We arrived back at 1800.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) Honda

Reason for use: David towed Opportunity back to the Hab.

Oil Added? No

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: 1

Notes on ATVs: None

HabCar used and why, where?

Atilla came from Grand Junction to MDRS.

Dean picked up water in Hanksville around 2000h to bring to the Green Hab.

CrewCar used and why, where? Not Used

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: EV suits normal, all batteries read > 12.5 volts

Summary of Hab operations: Checked SOC of electrical system as follows:

0730 SOC 100%, generator turned off by David

1300 SOC 100%, generator off

1900 SOC 66%, generator turned on by David

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAMM operations: Nothing to report

Summary of any observatory issues: Observation submitted March 28th for MDRS-14 has not gone through yet.

Summary of health and safety issues: None